OpenStudy (anonymous):

which of the binomials below is a factor of this trinomial x^2-x-12 a.x-3 b.x-6 c.x+3 d.x+6

OpenStudy (mathstudent55):

Do you know how to factor a binomial of this type?

OpenStudy (anonymous):

no

OpenStudy (mathstudent55):

I'll show you, then you'll be able to answer the question.

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

You have a trinomial of the type \(x^2 + ax + b\) a and b are just numbers, but notice that there is no number with the x^2.

OpenStudy (mathstudent55):

The first rule of factoring is always try to factor a common factor out of all terms.

OpenStudy (mathstudent55):

In your case, the coefficients (the number parts of the terms) are 1, -1, and -12. They don't have a common factor, so there is no common factor we can take out.

OpenStudy (mathstudent55):

That means we start factoring.

OpenStudy (anonymous):

okay how do you factor

OpenStudy (mathstudent55):

This is the procedure. Look at the last number. In this case it's -12. You need to find two numbers that multiply to -12. Also, these two numbers must add to the middle number, -1.

OpenStudy (mathstudent55):

Here are all the sets of two integers whose product is -12: -1 * 12 = -12 -12 * 1 = -12 -2 * 6 = -12 -6 * 2 = -12 -3 * 4 = -12 -4 * 3 = -12 Ok so far?

OpenStudy (anonymous):

Im thinking it is B?

OpenStudy (mathstudent55):

You mean B from my choices or B from the problem's choices?

OpenStudy (anonymous):

yes i understand.. thank you for helping. B from the problem

OpenStudy (mathstudent55):

No. It isn't. Follow me to the end of my explanation. It's almost done. Then you'll choose the correct choice.

OpenStudy (anonymous):

okay thanks

OpenStudy (mathstudent55):

You saw my listing of 6 different sets of integers that multiply to -12, right?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Now you need to choose the correct one. The correct one is the one in which the two numbers add to the middle coefficient. The middle coefficient is -1. Of all 6 choices I gave you, which one has two numbers that add to -1?

OpenStudy (anonymous):

-4+3

OpenStudy (mathstudent55):

-1 + 12 = 11 -12 + 1 = -11 -2 + 6 = 4 -6 + 2 = -4 -3 + 4 = 1 -4 + 3 = -1

OpenStudy (mathstudent55):

Exactly. That's correct. Now all you need to do is set up two sets of parentheses. ( )( )

OpenStudy (mathstudent55):

Then insert x in the left side of each one (x )(x ) Finally insert each number in the right side of each parentheses. (x - 4)(x + 3) We need to call 3, +3 to show an operation.

OpenStudy (anonymous):

What is a polynomial?¿

OpenStudy (mathstudent55):

That's it. That is how you factor a trinomial of the type \(x^2 + ax + b\) In summary: 1. Try to factor a common factor. 2. Find two numbers that multiply to b and add to a. 3. Call these numbers p and q. 4. The factorization is (x + p)(x + q)

OpenStudy (mathstudent55):

A ploynomial is a sum of terms. Each term is a number multiplied by 0 or more variables.

OpenStudy (anonymous):

Thank you for your help. Im doing apex thats why i needed help.So the answer will be C

OpenStudy (mathstudent55):

A monomial is one term. A binomial is two-term polynomial. A trinomial is a three-term polynomial.

OpenStudy (mathstudent55):

x + 3 is correct.

OpenStudy (anonymous):

I love you

OpenStudy (anonymous):

Thank you so much

OpenStudy (mathstudent55):

You're welcome. Keep studying and asking questions. That's how you learn.

OpenStudy (anonymous):

When a trinomial is factored as (x+m)(x+n) what is the sum of m and n?

OpenStudy (mathstudent55):

Look at my summary above. I'll copy it below to make it easier to see.

OpenStudy (mathstudent55):

Factor trinomial \(x^2 + ax + b\) Procedure: 1. Try to factor a common factor. 2. Find two numbers that multiply to b and add to a. 3. Call these numbers p and q. 4. The factorization is (x + p)(x + q)

OpenStudy (mathstudent55):

Notice that the numbers you call m and n, I called p and q.

OpenStudy (anonymous):

so it would be the coefficient term?

OpenStudy (mathstudent55):

m and n are numbers that add to the middle coefficient of the trinomial.

OpenStudy (mathstudent55):

The sum of m and n is the coefficient of the x-term.

OpenStudy (anonymous):

so it would be the constant term or the degree of the trinomial

OpenStudy (mathstudent55):

No. Not the constant term. The constant term is the last number that has no variable. Also, it's not the degree. The degree is the exponent of x, which in this case is 2.

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

Once again, the sum of m and n is the coefficient of the x-term.

OpenStudy (mathstudent55):

You're welcome.